OpenStudy (anonymous):

Here are Janet's steps for solving the system below by eliminating x. 2x + 3y = -1 5x + 8y = 0 Step One: Multiply the first equation by 5. 10x + 15y = -5 Step Two: Multiply the second equation by -2. -10x - 16y = 0 Step Three: Add the two equations. y = -5 Step Four: Substitute y back into the first equation. 2x + 3(-5) = -1 Step Five: Solve for x. 2x - 15 = -1 2x = 14 x = 7 Is she correct? Justify your answer. If incorrect, describe the error.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

haha,as soon as i mention you

OpenStudy (anonymous):

Look at step 3, do you notice anything?

OpenStudy (anonymous):

shouldnt it be y=5

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

yeshh

OpenStudy (anonymous):

So you can see from step 3 and onwards, the whole problem will be wrong then.

OpenStudy (texaschic101):

2x + 3y = -1 --->(5) 5x + 8y = 0 --->(-2) --------------------- 10x + 15y = -5 (result f multiplying by 5) -10x - 16y = 0 (result of multiplying by -2) --------------------add -y = -5 y = 5 (This is where Janet made her mistake) now sub 5 in for y in either equation to find x 2x + 3y = -1 2x + 3(5) = -1 2x + 15 = -1 2x = -15 - 1 2x = -16 x = -8 check.. 2x + 3y = -1 2(-8) + 3(5) = -1 -16 + 15 = -1 -1 = -1 (correct) x = -8 and y = 5 and Janet is incorrect

OpenStudy (anonymous):

thank you again miss texas, youre a life saver

OpenStudy (texaschic101):

well...I don't know about a lifesaver...but thank you :)

OpenStudy (anonymous):

pretty close